# The Unpaid Toll: Quantifying the Public Health Impact of AI

### v1

@@ Appendix D Health-informed AI
Table 7: Correlation analysis of marginal carbon emissions and health impacts for Google’s U.S. data center locations between October 1, 2023, and September 30, 2024 [71]. According to the region classification of WattTime [95], the two data centers in Storey County, NV, and Henderson, NV, belong to the same power grid region, and so do those in Jackson County, AL, and Montgomery County, TN.
D.2 Location-dependent Public Health Impact
We now show the location-dependent public health impacts of two technology companies based on Google’s and Meta’s U.S. data center locations in 2023, excluding their leased colocation data centers whose locations are proprietary. Due to the lack of information about the per-data center electricity consumption, we uniformly distribute Google’s North America electricity consumption over its U.S. data center locations based on Google’s latest sustainability report [10]. Meta discloses its per-location electricity usage [45]. We consider criteria air pollutants without accounting for renewable energy credits these two companies apply to offset their grid electricity consumption (see “Location-based emission” in Appendix A.2). As a consequence, although we consider the U.S. data center locations of Google and Meta, our results should not be interpreted as a quantitative evaluation of these two specific companies’ actual public health impacts. We also emphasize that our goal is to highlight the locational dependency of public health impacts and to motivate the need for health-informed siting of data centers. In our results, we refer to Google and Meta as Company A and Company B, respectively, to avoid potential misunderstandings.
We first see from Table 8 that while the two companies have different public health costs due to their different electricity consumption, their health-to-electricity cost ratios are similar at the national level. Nonetheless, we notice from Fig. 12 that the two companies have significant differences in terms of the per-household health cost distribution and most-affected counties. This is primarily due to the two companies’ different data center locations, and highlights the locational dependency of public health impacts. That is, unlike carbon emissions that have a similar effect on climate change regardless of the emission source locations, the public health impact of criteria air pollutants heavily depends on the location of the emission source. Thus, technology companies should account for public health impacts when deciding where they build data centers, where they get electricity for their data centers, and where they install renewables in order to best mitigate the adverse health effects while promoting equity.
Table 8: The public health costs based on two technology companies’ U.S. data center electricity consumption in 2023.
Company (Attribution) Electricity (TWh) Electricity Cost (billion \$) Health Cost (billion \$) % of Electricity Cost Per-Household Health Cost (\$)
A (Average) 18.5 1.4 0.63 (0.47, 0.78) 45% 4.5 (3.4, 5.5)
A (Marginal) 18.5 1.4 0.97 (0.75, 1.20) 70% 6.9 (5.3, 8.6)
B (Average) 10.6 0.8 0.38 (0.29, 0.48) 51% 2.7 (2.0, 3.4)
B (Marginal) 10.6 0.8 0.53 (0.41, 0.66) 71% 3.8 (2.9, 4.7)
(a) Per-household health cost (Company A)
(b) Per-household health cost (Company B)
(c) CDF of per-household health cost (Company A)
(d) CDF of per-household health cost (Company B)
State County Per-household Health Cost (\$) County-to-nation Income Ratio TX Marion 33.8 (27.2, 40.4) 0.64 VA Mecklenburg 23.7 (19.3, 28.1) 0.68 VA Halifax 23.5 (18.8, 28.1) 0.65 NC Person 23.5 (19.4, 27.6) 0.81 VA Martinsville City 22.5 (18.6, 26.5) 0.52 VA Danville City 21.7 (17.3, 26.1) 0.55 TX Cass 21.2 (16.5, 25.8) 0.72 GA Pickens 21.1 (17.3, 24.9) 0.97 WV Marion 20.5 (16.3, 24.6) 0.80 VA Henry 20.4 (15.9, 24.8) 0.58
(e) Top-10 counties by per-household health cost (Company A)
State County Per-household Health Cost (\$) County-to-nation Income Ratio TX Marion 21.1 (17.0, 25.3) 0.64 TX Cass 13.3 (10.4, 16.3) 0.72 WV Marion 12.4 (9.9, 15.0) 0.80 GA Pickens 12.4 (10.1, 14.6) 0.97 WV Marshall 12.1 (9.2, 14.9) 0.77 WV Mason 12.0 (9.4, 14.6) 0.71 TX Gregg 12.0 (9.6, 14.5) 0.85 TX Harrison 12.0 (9.7, 14.3) 0.84 TX Morris 12.0 (9.3, 14.6) 0.69 OH Gallia 11.9 (8.9, 14.9) 0.74
(f) Top-10 counties by per-household health cost (Company B)
Figure 12: The county-level per-household health cost of two companies in 2023. The income data is based on the 2018-2022 American Community Survey 5-year estimates provided by [83]. Average attribution.

### v2

@@ 6.1 Opportunities for Health-Informed Computing
Table 3: Correlation analysis of marginal carbon emissions and health impacts for a technology company’s U.S. data center locations between October 1, 2023, and September 30, 2024 [77]. According to the region classification of WattTime [98], the two data centers in Storey County, NV, and Henderson, NV, belong to the same power grid region, and so do those in Jackson County, AL, and Montgomery County, TN.
In addition to analysis for all U.S. regions, we turn to specific regions where a large technology company builds its U.S. data centers. We present the results Table 3, further confirming that carbon intensities and health impacts are not always aligned and that health impacts vary more significantly than carbon intensities in almost all the locations.
We further analyze the Pearson correlation coefficients between hourly marginal health prices and carbon emission rates throughout 2023 for the U.S. regions that have complete health and carbon data provided by WattTime [77]. The CDF of the correlation coefficients is shown in Figure 9a. We see that nearly 70% of the regions have a weak or moderate correlation, with a carbon-health correlation coefficient of less than 0.60. This implies that despite having fossil fuels as the common source, health costs and carbon emissions are different and can exhibit trade-offs. Moreover, due to their additional dependence on population distribution and meteorological conditions, health prices demonstrate more pronounced temporal fluctuations than carbon emissions.
These findings highlight that leveraging spatiotemporal variations in a health-aware manner can substantially reduce the public health costs of data center operations. Moreover, the observed distinctions between health impacts and carbon emissions suggest the need to optimize data center decisions by explicitly accounting for and exploiting the spatiotemporal heterogeneity of health impacts.
6.2 Benefits of Health-Informed Computing
To improve system performance and reliability, technology companies typically operate data centers over a variety of geographically distributed regions and dynamically distribute computing workloads through a process known as geographical load balancing (GLB). Here, we leverage the unique spatial load flexibility of geographically distributed data centers to demonstrate the benefits of health-informed computing as a proof of concept.
Health-Informed GLB. Specifically, we study health-informed GLB (HI-GLB) to address the public health burden of data center operation. We consider a discrete-time model of duration $T$ and measure the workloads in terms of their energy demand. For the sake of examining the impact of spatial flexibility, we assume that the energy loads (i.e., workloads) can be flexibly distributed across a set of $N$ data centers denoted as $\mathcal{N}=\{1,2,\ldots,N\}$. In each time $t\in\{1,2,\ldots,T\}$, the total energy demand is $M_{t}$, and $w_{i,t}$ represents the load assigned to data center $i$. We use $l_{i}$ to represent the default load capacity of data center $i$, and introduce a slackness parameter $\lambda\geq 1$ to represent the ability of each data center to accept loads in excess of its default capacity. Thus, the greater the value of $\lambda$, the more spatial flexibility the operator has. In each time $t$, we use $p_{i,t}^{e}$ and $p_{i,t}^{h}$ to denote the electricity price and health price at data center $i$, respectively.
The health price $p_{i,t}^{h}$ depends on the emission source of criteria air pollutants (e.g., power plants’ emission rates and their locations), air pollutant dispersion, and estimates of adverse health outcomes and resulting costs attributed to the increased air pollutant concentration in each region [22]. Thus, the health price quantifies the ultimate health burden imposed on affected populations and is measured in terms of economic costs for each unit of electricity consumption. It varies over time due to fluctuations in the grid’s generation mix and changing meteorological conditions. Third-party organizations such as WattTime [98] provide real-time estimates of the marginal health price of electricity across 114 power balancing regions in the U.S., while the EPA [100] reports annualized average health prices for electricity in 14 broader regions nationwide.
Our goal is to minimize the sum of electricity costs $\sum_{t=1}^{T}\sum_{i=1}^{N}p_{i,t}^{e}\cdot w_{i,t}$ and health costs $\sum_{t=1}^{T}\sum_{i=1}^{N}p_{i,t}^{h}\cdot w_{i,t}$ across all regions. Thus, HI-GLB is formulated as follows:
$\min_{{w}}$ $\qquad\sum_{t=1}^{T}\sum_{i=1}^{N}p_{i,t}^{e}\cdot w_{i,t}+\sum_{t=1}^{T}\sum_{i=1}^{N}p_{i,t}^{h}\cdot w_{i,t}$ (4a)
$\mathrm{s.t.}$ $\qquad\sum_{i=1}^{N}w_{i,t}=M_{t},\quad\forall t\in\mathcal{T}$ (4b)
$\qquad 0\leq w_{i,t}\leq\lambda\cdot l_{i},\quad\forall i\in\mathcal{N},\forall t\in\mathcal{T}$ (4c)
where the constraint (4b) means that all loads must be dispatched to a data center (with no loads dropped) and the constraint (4c) encodes the maximum workload capacity of each data center.
Our formulation can be easily extended to incorporate additional considerations such as load dispatching distance constraints and other metrics such as carbon emissions. Additionally, it can also include long-term, per-region health impact constraints, rather than focusing solely on national-level total health costs. For the sake of clarity, we set these extensions aside to focus on the novel metric of health cost for data center resource management.
[…]
For each data center location, we use the health price provided by WattTime [77] and the industry electricity price from the EIA [102]. To quantify the carbon emissions of each location, we consider the marginal carbon emissions rate provided by WattTime [77]. We specifically use marginal carbon emissions for a fair comparison with the marginal health price, and since this metric often better reflects the actual impact on system-wide carbon emissions resulting from demand-side load shifting [99]. More details can be found in Appendix B.1.
Table 4: Comparison between health-informed and carbon-aware GLB ($\lambda=1.5$)
Metric Baseline Carbon-Aware GLB Health-Informed GLB
$p^{c}=$ \$0/ton $p^{c}=$ \$5/ton $p^{c}=$ \$200/ton $p^{c}=$ $\infty$/ton HI-GLB ($p_{i,t}^{e}=0$) HI-GLB
Health (Million \$) 393.23 416.29 (5.86%) 416.29 (5.86%) 383.32 (-2.52%) 404.26 (2.80%) 289.67 (-26.34%) 291.94 (-25.76%)
Energy (Million \$) 756.50 714.99 (-5.49%) 714.99 (-5.49%) 734.77 (-2.87%) 765.68 (1.21%) 741.49 (-1.98%) 733.66 (-3.02%)
Carbon (Million Ton) 6.60 6.67 (1.02%) 6.67 (1.02%) 6.20 (-6.01%) 6.12 (-7.23%) 6.54 (-0.89%) 6.51 (-1.38%)
Figure 9: Correlation analysis. (a) CDF of correlation coefficients between hourly health prices and marginal carbon emission rates for all the U.S. regions; (b) Scatter plot of health price and marginal carbon emission rate (annual average in 2023) across Meta’s U.S. data center locations.
Health-Informed vs. Carbon-Aware GLB. We show results comparing GLB with health and carbon costs in Table 4. The results highlight that optimizing solely for environmental metrics like carbon does not effectively mitigate health impacts. Concretely, as the carbon price varies, the health cost ranges between \$380 million and \$420 million, whereas the baseline health cost is approximately \$393 million. When $p_{c}=\mathdollar 200$/ton, the health cost decreases by 2.52% relative to the baseline. However, for all other carbon prices, the health cost increases by using carbon-aware GLB. Among these, the algorithm that focuses solely on minimizing carbon ($p_{c}=\infty$) leads to a 2.80% increase in health cost compared to the baseline. This stands in sharp contrast to HI-GLB, which achieves a substantial $\sim$26% reduction in health cost compared to the baseline, demonstrating that environmental metrics such as carbon are insufficient for mitigating health impacts unless public health considerations are explicitly integrated into the optimization process. Moreover, compared to the pure health-informed GLB that sets a zero electricity price ($p^{e}=0$), HI-GLB only slightly increases the health cost from \$290 million to \$292 million, further demonstrating that considering the health price can effectively drive healthier GLB decisions while maintaining low cost.
At the same time, the pure carbon-aware GLB algorithm achieves a maximum reduction in carbon reduction of 7.23% while increasing the health burden and electricity cost relative to the baseline. In contrast, HI-GLB achieves 26% health cost reduction, 3% electricity cost saving, and more than 1% carbon reduction. This highlights the necessity of adopting health-informed algorithms for health burden reduction while offering co-benefits of cost saving and carbon reduction.
We also vary the capacity slackness $\lambda$ to change the spatial flexibility and obtain similar insights. The full details are available in Appendix B.2.
Spatial correlation analysis. To further explain the results, we analyze the correlation between health costs and carbon emissions across the 13 regions where Meta operates its U.S. data centers. Figure 9b presents a scatter plot illustrating the correlation between the annual average carbon emission rate and health price across various regions. The Pearson correlation coefficient between different locations’ health prices and carbon emissions is approximately -0.35, indicating a negative relationship. This suggests a potential conflict between efforts to optimize carbon emissions and those aimed at improving health outcomes via GLB. Furthermore, as highlighted by the spatial patterns in Figure 9b, the health prices exhibit a significantly higher degree of spatial variability compared to carbon emissions across regions. This discrepancy further reinforces the point that focusing solely on optimizing environmental factors, such as carbon emissions, may not effectively reduce health costs, underscoring the importance of integrating health-informed optimization strategies to achieve more comprehensive benefits.
7 Our Recommendations
We provide additional recommendations to address the growing public health impact of data centers.
Recommendation 1: Standardization of Reporting Protocols
Despite their immediate and tangible impacts on public health, criteria air pollutants have been entirely overlooked in AI model cards and sustainability reports published by technology companies [37, 36, 10]. The absence of such critical information adds substantial challenges to accurately identifying specific AI data centers as a key root cause of public health burdens and could potentially pose hidden risks to public health. To enhance transparency and lay the foundation for truly responsible AI, we recommend standardization of reporting protocols for criteria air pollutants and the public health impacts across different regions. Concretely, criteria air pollutants can be categorized into three different scopes (Section 3), and reported following the greenhouse gas protocol widely adopted by technology companies [37, 10, 78].
Recommendation 2: Attention to All
Recommendation 3: Promoting Public Health Equity
The public health impact of AI is highly unevenly distributed across different counties and communities, disproportionately affecting certain (often low-income) communities [103, 31]. For example, as shown in Table 6c, all the top-10 most impacted counties in the U.S. have lower median household incomes than the national median value. The ratio of the highest county-level per-household health cost to the lowest cost is approximately 200. Therefore, it is imperative to address the substantial health impact disparities across communities.
8 Conclusion
In this paper, we quantify and address the overlooked public health impact of data centers. We introduce a principled methodology to model these lifecycle pollutant emissions and quantify their associated public health impacts. Our findings suggest that the total annual public health burden of U.S. data centers could exceed \$20 billion by 2028, approaching or even surpassing the impacts of on-road vehicle emissions in California. Importantly, these health costs are not evenly distributed: disadvantaged communities bear a disproportionate share, with per-household impacts potentially up to 200 times higher than in less-affected areas. This highlights the need for targeted mitigation strategies. To this end, we propose health-informed computing, a framework that explicitly incorporates public health risk as a key metric when scheduling data center workloads, enabling more informed and equitable operational decisions.
More broadly, we recommend the adoption of standardized reporting protocols for the public health costs of data centers, alongside policies that ensure attention to all impacted communities, thereby supporting responsible, sustainable, and inclusive deployment of AI infrastructure.
Acknowledgement
Appendix
Appendix A Modeling Details
We describe the evaluation methodology used for our empirical analysis. We use the latest COBRA (Desktop v5.1, as of October 2024) provided by the U.S. EPA [39] to study the public health impact of U.S. data centers in both 2019-2023 and 2028. While COBRA uses a reduced-complexity air quality dispersion model based on a source-receptor matrix for rapid evaluation, its accuracy has been validated and the same or similar model has been commonly adopted in the literature for large-area air quality and health impact analysis [85, 87, 82, 88]. We consider county-level air pollutant dispersion throughout the contiguous U.S., which is the area currently supported by COBRA [39]. Note that cities considered county-equivalents for census purposes are also referred to as “counties” in COBRA. Throughout the paper, we use “county” without further specification.
COBRA provides data for county-level population, health incidence, valuation, and baseline emissions for 2016, 2023, and 2028 [39]. For the data from 2019 to 2022, we use linear interpolation as recommended by the EPA’s COBRA team.
We show in Table 5 and Table 6 the total baseline emissions of air pollutants for electricity generation and on-road traffic provided by COBRA [39]. By reducing a state’s on-road emissions to zero in COBRA, we obtain the corresponding public health cost in that state.
Table 5: U.S. electricity generation baseline emissions from 2016 to 2028
Year Electricity Generation Emission (Metric Ton)
NOx SO2 PM2.5 VOC
2016 1100575.41 1369417.44 111604.62 30250.76
2023 711746.94 717409.25 110878.22 34311.54
2028 695495.34 733437.11 110279.40 34446.71
Table 6: U.S. and California on-road baseline emissions from 2016 to 2028
NOx SO2 PM2.5 VOC NOx SO2 PM2.5 VOC
On-road emissions are categorized as the “Highway Vehicles” sector in COBRA and include both tailpipe exhaust and tire and brake wear. Thus, following the EPA and U.S. Department of Transportation classification [22, 104], PM2.5 resulting from road dust is not counted as emissions of highway vehicles in our study. If the PM2.5 from paved road dust (categorized as “Miscellaneous $\to$ Other Fugitive Dust $\to$ Paved Roads” in COBRA) is considered, California is still projected to have the highest state-wide public health cost of on-road vehicles among all the U.S. states.
Electricity price. When estimating the electricity cost for data centers in 2023 and 2038, we use the state-level average price for industrial users in [102]. The projected U.S. nominal electricity price for industrial users remains nearly the same from 2023 to 2030 (24.96 \$/MMBtu in 2023 vs. 23.04 \$/MMBTu in 2030) in the baseline case per the EIA’s Energy Outlook 2023 [105]. Thus, our estimated health-to-electricity cost ratio will be even higher if we further adjust inflation.
A.1 Public Health Impact of Backup Generators in Virginia
The scope-1 emission information for data centers in other states is not always publicly available. Thus, when estimating scope-1 emissions for data centers in another state, we apply the emission rate (tons/MWh) derived from Virginia’s data and multiply it by the energy consumption of data centers in that states. Although this approach may introduce some estimation errors, the impact is expected to be limited because scope-2 health costs are substantially more dominant.
A.2 Data Centers’ Scope-2 Public Health Impact
The locations of emission sources depend on the power plants supplying electricity to data centers. To evaluate the public health impacts of U.S. data centers, we focus on average attribution method, which is also the standard methodology of carbon emission accounting used by technology companies in their sustainability reports [10, 78, 37].
We first calculate the total data center electricity consumption $e_{DC}$ and the overall electricity consumption (including non-data center loads) $e_{Total}$ within each electricity region. The U.S. electricity grid is divided into 14 regions following the AVoided Emissions and geneRation Tool (AVERT, the latest version v4.3 as of October 2024) provided by the EPA [76]. We use the state-level data center electricity consumption distribution for 2023 provided by EPRI [5], scale it by the U.S. total data center electricity consumption in 2019-2023 and for the 2028 projection based on data provided by [4], and then distribute state-level electricity consumption to relevant electricity regions following the state-to-region electricity apportionment used by AVERT.
A.3 Public Health Impact of a Semiconductor Facility
Although semiconductor manufacturing facilities are subject to air quality regulations [74], they still pose significant risks, affecting populations across large regions. Maricopa County, AZ, has been an EPA-designated non-attainment area for several years due to its failures to meet federal air quality standards [106]. The establishment of multiple semiconductor facilities in such areas could further exacerbate air quality issues.
A.4 Energy Consumption for Training a Generative AI Model
We consider Llama-3.1 as an example generative AI model. According to the model card [36], the training process of Llama-3.1 (including 8B, 70B, and 405B) utilizes a cumulative of 39.3 million GPU hours of computation on H100-80GB hardware, and each GPU has a thermal design power of 700 watts. Considering Meta’s 2023 PUE of 1.08 [37] and excluding the non-GPU overhead for servers, we estimate the total training energy consumption as approximately 30 GWh. Our estimation method follows Meta’s guideline [36] and is conservative, as it excludes the substantial non-GPU energy overheads (e.g., CPUs) associated with server operations.
A.5 Average Emission for Each LA-NYC Round Trip by Car
A.6 State-wide Electricity Consumption by U.S. Data Centers in 2023
We show in Fig. 10 the state-wide data center electricity consumption in 2023 [5]. It can be seen that Virginia, Texas and California have the highest data center electricity consumption in 2023. The total national electricity consumption reported by EPRI is slightly lower than the values in [4], and we scale it up accordingly in our calculations to ensure consistency.
(a) Electricity consumption map
State Electricity Consumption (TWh) State Electricity Consumption (TWh) State Electricity Consumption (TWh) VA 33.85 OH 2.36 ID 0.15 TX 21.81 SC 2.02 WI 0.15 CA 9.33 WY 1.86 MD 0.10 IL 7.45 KY 1.62 LA 0.08 OR 6.41 CO 1.51 SD 0.07 AZ 6.25 AL 1.49 ME 0.03 IA 6.19 FL 1.38 NH 0.02 GA 6.18 TN 1.33 RI 0.02 WA 5.17 OK 1.23 KS << 0.01 PA 4.59 MA 1.06 AR << 0.01 NY 4.07 MO 0.97 DE << 0.01 NJ 4.04 MN 0.82 DC << 0.01 NE 3.96 MT 0.58 MS << 0.01 ND 3.92 MI 0.53 VT << 0.01 NV 3.42 NM 0.40 WV << 0.01 NC 2.67 CT 0.26 UT 2.56 IN 0.19
(b) Electricity consumption by state (descending order)
Appendix B Additional Results for Health-Informed GLB
B.1 Details of the experiment setup
We use Meta’s electricity consumption for each U.S. data center location in 2023 [37] for our experiments. Table 7 summarizes the baseline annual energy load $W_{i}$ for each data center $i$. Since data centers are mostly stable loads in practice, the hourly workload for each location is calculated as $l_{i}=\frac{W_{i}}{T}$, where $T$ represents the total number of hours in the study period. The total hourly workload is then computed as $M_{t}=\sum_{i=1}^{N}l_{i}$.
We use the annual average industrial electricity prices in different states provided by the EIA [109]. The health price $p_{i,t}^{h}$ (\$/MWh) and carbon emission rate $r_{i,t}^{c}$ (ton/MWh) are based on data provided by WattTime [77]. WattTime divides the U.S. into more than 100 regions and provides marginal health prices and carbon emission rates for each region. These values are updated every 5 minutes.
In our study, we use the WattTime data from 0:00 on January 1, 2023, to 23:55 on December 31, 2023. The actual time interval for our experiment is 1 hour, where the 5-minute data points are averaged hourly to compute $p_{i,t}^{h}$ and $r_{i,t}^{c}$, $\forall t=\{1,2,\ldots,T\}$, with $T=8760$.
B.2 Sensitivity analysis of capacity slackness $\lambda$
The parameter $\lambda\geq 1$ represents the capacity slackness to accept additional loads: The greater $\lambda$, the more spatial flexibility there is. Here, we vary $\lambda$ and report the results for $\lambda=1.2$ and $\lambda=2.0$ in Tables 8 and 9, respectively. Similar to Table 4, the potential reduction in health costs is significant. An increase in $\lambda$ corresponds to an expansion in the spatial flexibility, leading to a larger reduction in health cost. Specifically, when $\lambda=2.0$, the health-informed algorithm achieves a reduction in health cost of over 40% compared to the baseline. In contrast, under $\lambda=2.0$, the pure carbon-aware algorithm, which focuses exclusively on optimizing carbon emissions with an infinite carbon price $p^{c}=\infty$, can achieve a reduction in carbon emissions of approximately 12% relative to the baseline. This comparison highlights the great potential for health cost reduction and reinforces the importance of designing HI-GLB.
Table 7: Information about Meta’s U.S. data center locations in 2023
Location Energy (MWh) Electricity Price $p^{e}$ (\$/MWh) Health Price $p^{h}$ (\$/MWh) Carbon Intensity $r^{c}$(ton/MWh)
Huntsville, AL 614198 71.0 49.49 0.59
Stanton Springs, GA 968565 68.8 19.35 0.58
DeKalb, IL 138965 82.0 51.72 0.58
Altoona, IA 1243306 69.1 57.86 0.65
Sarpy, NE 1148091 76.3 36.72 0.56
Los Lunas, NM 1110100 57.5 18.96 0.68
Forest City, NC 507068 71.5 54.64 0.66
New Albany, OH 793063 70.3 54.32 0.61
Prineville, OR 1375321 75.2 13.31 0.65
Gallatin, TN 116520 62.3 49.49 0.59
Fort Worth, TX 1029570 66.0 44.56 0.51
Eagle Mountain, UT 787740 69.9 21.93 0.74
Henrico, VA 805061 89.2 54.37 0.61
Table 8: Comparison between health-informed and carbon-aware GLB ($\lambda=1.2$)
Metric Baseline Carbon-Aware GLB Health-Informed GLB
$p^{c}=$ \$0/ton $p^{c}=$ \$5/ton $p^{c}=$ \$200/ton $p^{c}=$ $\infty$/ton HI-GLB ($p_{i,t}^{e}=0$) HI-GLB
Health (Million \$) 393.23 374.20 (-4.84%) 374.44 (-4.78%) 395.28 (0.52%) 404.72 (2.92%) 345.26 (-12.20%) 349.29 (-11.17%)
Energy (Million \$) 756.50 732.06 (-3.23%) 732.30 (-3.20%) 747.88 (-1.14%) 761.20 (0.62%) 752.26 (-0.56%) 738.49 (-2.38%)
Carbon (Million Ton) 6.60 6.68 (1.22%) 6.56 (-0.60%) 6.40 (-3.04%) 6.37 (-3.54%) 6.54 (-0.87%) 6.57 (-0.44%)
Table 9: Comparison between health-informed and carbon-aware GLB ($\lambda=2.0$)
Metric Baseline Carbon-Aware GLB Health-Informed GLB
$p^{c}=$ \$0/ton $p^{c}=$ \$5/ton $p^{c}=$ \$200/ton $p^{c}=$ $\infty$/ton HI-GLB ($p_{i,t}^{e}=0$) HI-GLB
Health Cost (Million \$) 393.23 368.10 (-6.39%) 409.67 (4.18%) 374.82 (-4.68%) 408.92 (3.99%) 221.38 (-43.70%) 223.96 (-43.05%)
Energy Cost (Million \$) 756.50 702.16 (-7.18%) 702.69 (-7.11%) 723.42 (-4.37%) 767.59 (1.47%) 734.29 (-2.94%) 728.19 (-3.74%)
Carbon (Million Ton) 6.60 6.72 (1.73%) 6.52 (-1.17%) 5.95 (-9.91%) 5.84 (-11.49%) 6.61 (0.19%) 6.55 (-0.71%)